OpenStudy (anonymous):

derivative [-sqrt(x^2+4)/2x] - (1/4) ln[[(2+sqrt(x^2+4)]/x] How to solve it?

OpenStudy (anonymous):

to do a problem like that you need to better express it. Hard to follow along. Try demonstranda.com and see who is online. Its new, but they have a better solver in MathMatix

OpenStudy (anonymous):

\[\frac {-\sqrt{x^{2}+4}}{2x} - \frac {1}{4} \ln(\frac {2+\sqrt{x^{2}+4}}{x})\]

OpenStudy (anonymous):

this is how it look likes, what should i do first

OpenStudy (anonymous):

chain and quotient rules